# MobilityGPT: Enhanced Human Mobility Modeling with a GPT model

### v2

@@ B.1 Benchmark models
  • SeqGAN (Yu et al., 2017): SeqGAN is a state-of-the-art sequential generative model that trains trajectories in discrete sequences with RL policy gradient methods using Monte Carlo sampling.
  • LSTM-TrajGAN (Rao et al., 2020): This is a tabular LSTM model by learning the normalized GPS location divergences with respect to the center of the dataset.
  • TS-TrajGen (Jiang et al., 2023): TS-TrajGen is a method based on GANs that incorporates a modified $A*$ path search algorithm. This adjustment ensures spatial continuity in the generated data. Additionally, the method includes topological constraints to align the generated trajectories with existing road networks, ensuring road matching. TS-TrajGen requires fixed OD pairs as inputs. We use OD pairs from the real training set as inputs during inference.
  • DiffTraj (Zhu et al., 2024): is the most recent and promising generative model for mobility trajectories. DiffTraj employs a diffusion model in the U-Net structure for generating GPS trajectories.
B.2 Evalaution metrics
To evaluate the performance of the generative models, we need a proper comparison metric with a wide range of mobility characteristics that quantifies the quality of generated trajectories given the real trajectories. In this project, we employed four well-known comparison metrics from literature (Hasan et al., 2013; Gursoy et al., 2018) in addition to our proposed gravity and connectivity utility metrics. These metrics have been widely used for the utility of synthetic mobility data generation models (Haydari et al., 2022; Zhu et al., 2024).
@@ B.2 Evalaution metrics
Apart from the query error, the main utility metrics inspect the distributional similarities for synthetic trajectories with respect to real trajectories that train the generative model. Jensen-Shannon divergence (JSD) is a well-known similarity metric mainly used for measuring the similarity of two probability distributions (Lin, 1991). We employ JSD for several human mobility characteristics.
@@
Connectivity: The connectivity metric quantifies the percentage of trajectories that are fully connected from origin to destination, owing to the RCM that samples the next road link conditioned on the connected road links, thereby capturing the underlying road network topology and connectivity patterns.
B.3 Model Details
MobilityGPT utilizes a minimal version of the standard decoder-only transformer architecture (Radford et al., 2019), built upon the minimal GPT implementation from (Karpathy, ). Table 3 summarizes the important model and training parameters employed in this work. Since the Porto and Beijing datasets exhibit different trajectory lengths and characteristics, we utilize distinct block sizes for training and maximum trajectory lengths for synthetic trajectory generation. However, we employ the same set of hyperparameters across both datasets for all training stages, including the reward model and PPO training.
Table 3: Model and training parameters utilized in MobilityGPT
@@ B.3 Model Details
Appendix C Additional Experiments
C.1 Ablation Study on Beijing Dataset
An additional ablation study on the Beijing dataset, shown in Table C.1, investigated the impact of MobilityGPT’s key components. Similar to the Porto dataset, the ablation study on the Beijing dataset exhibits the synergistic effects of MobilityGPT’s novel components – RCM, Gravity Model, and RLTF fine-tuning – in accurately modeling diverse mobility characteristics, boosting performance on key metrics like OD and Trip Length while the connectivity priors ensure accurate capture of network-level patterns.
Table 4: Ablation on Beijing dataset study for different components of MobilityGPT.
       RCM Gravity Model RLTF OD Trip Length Connectivity
0.102 0.201 0.330
0.103 0.151 0.367
0.093 0.145 1.0
0.099 0.131 1.0
0.099 0.123 1.0
C.2 Effect of temperature on MobilityGPT inferencing
The temperature parameter in LLMs, including MobilityGPT, influences the randomness in generating outputs. At lower temperatures, the model behavior is more deterministic, often selecting the most likely next token. As the temperature increases, the model introduces greater randomness, leading to a diverse array of outputs.
Interestingly, as illustrated in Figure 5, MobilityGPT demonstrates a clear trend: its performance tends to increase with higher temperature settings. This observation suggests that mobility data generation can benefit from the randomness of GPT-like models.
Figure 5: Performance of MobilityGPT under different temperature settings.
C.3 Tokenizer Modeling
In the development of the MobilityGPT tokenizer model, we explore the use of specific tokens to mark the start (<BOT>) and end (<EOT>) of trajectories. Our aim is to assess how these tokens affect the model’s performance in human mobility modeling, particularly for sequential trajectory data. Our findings reveal a key insight: using the <EOT> token alone significantly improves performance. This result indicates that emphasizing the end of a trajectory is more beneficial than marking its beginning with a separate token for our mobility modeling task. The effectiveness of the <EOT> token emphasizes the importance of nuanced data representation and the need for task-specific tokenization strategies in sequential data generation with transformer models. An ablation study, detailed in Table 5, compares single and dual-token approaches using various metrics. This study shows that using a single token type outperforms the dual-token method in the MobilityGPT workflow, highlighting the advantages of customized tokenization.
Table 5: Ablation study on the effectness of <EOT>.
       <BOT> <EOT> OD Trip Length Radius Gravity
0.110 0.151 0.128 0.235
0.114 0.124 0.107 0.225
Appendix D Computational Demand
Evaluating computational demand is essential for ensuring the practical feasibility, scalability, and environmental sustainability of generative models. To facilitate a fair comparison, we trained all benchmark models and MobilityGPT on the same computational resources, utilizing an NVIDIA TITAN RTX GPU with 24GB memory, and the results are depicted in Table 6. MobilityGPT balances model size and efficiency: $5.5M$ params, lower than TS-TrajGen ($23.2M$ $7.6M$) but higher than LSTM-TrajGAN ($100k$). Its $\sim$5hr runtime outperforms TS-TrajGen ($>200hr$) and DiffTraj ($48-60hr$) by an order of magnitude, despite LSTM-TrajGAN’s $30min-1hr$ efficiency at the cost of poor predictive performance.
Table 6: Computational performance of different models vs MobilityGPT
  N_param Total running time
Model BJ Porto BJ Porto
SeqGAN 2.6M 800 $\sim$ 12hr $\sim$ 5hr
LSTM-TrajGAN 100k 100k $\sim$ 30mins $\sim$ 1hr
TS-TrajGen 23.2M 7.6M $>$ 200hr $>$ 200hr
DiffTraj 15.7M 15.7M $\sim$ 48hr $\sim$ 60hr
MobilityGPT 5.5M 5.5M $\sim$ 5hr $\sim$ 5hr
MobilityGPT achieves a compelling balance between model capacity and scalability through its moderate size and computational efficiency. Despite having fewer parameters than some benchmarks, MobilityGPT achieves state-of-the-art performance across datasets, requiring substantially less time and resources. Unlike LSTM-TrajGAN, which trades off predictive accuracy for efficiency, MobilityGPT demonstrates the feasibility of combining high performance and practical efficiency within a single model. This makes MobilityGPT a promising choice for practical deployments where both predictive power and computational feasibility are essential.

### v3

@@ V-B Benchmark models
  • Random Walk: This baseline randomly selects a starting point on the road network and performs a random walk guided by visitation counts of the road segments (locations). It operates solely on the road graph structure without learning from the data distribution.
  • Mobility Markov Chain (MMC) [gambs2012next]: This model treats road segments as discrete states and constructs a transition probability matrix to capture first-order dependencies between segments. The next segment is sampled based on the learned transition probabilities.
  • SeqGAN [yu2017seqgan]: SeqGAN is a state-of-the-art sequential generative model that trains trajectories in discrete sequences with RL policy gradient methods using Monte Carlo sampling.
  • LSTM-TrajGAN [rao2020lstm]: This is a tabular LSTM model by learning the normalized GPS location divergences with respect to the center of the dataset.
  • TS-TrajGen [jiang2023continuous]: TS-TrajGen is a method based on GANs that incorporates a modified AA* path search algorithm. This adjustment ensures spatial continuity in the generated data. Additionally, the method includes topological constraints to align the generated trajectories with existing road networks, ensuring road matching. TS-TrajGen requires fixed OD pairs as inputs. We use OD pairs from the real training set as inputs during inference.
  • DiffTraj [zhu2024difftraj]: is the most recent and promising generative model for mobility trajectories. DiffTraj employs a diffusion model in the U-Net structure for generating GPS trajectories.
V-C Evaluation metrics
To evaluate the performance of the generative models, we require robust comparison metrics that capture a broad spectrum of mobility characteristics and effectively quantify the quality of the generated trajectories relative to real ones. The selected metrics assess whether the synthetic trajectories preserve both local transition patterns—i.e., movements between adjacent road segments—and global route structures that reflect realistic and rational human mobility behavior in urban environments. In this study, we adopted four widely used evaluation metrics, along with our proposed gravity and connectivity utility metrics, as introduced in prior works [hasan2013spatiotemporal, gursoy2018utility]. These metrics have become standard tools for assessing the utility of synthetic mobility data generation methods [haydari2022differentially, zhu2024difftraj]. Given that MobilityGPT introduces randomness and can alter trajectory patterns, direct one-to-one trajectory comparisons are not applicable in such generative settings [kapp2023generative].
@@ V-C Evaluation metrics
Apart from the query error, the main utility metrics inspect the distributional similarities for synthetic trajectories with respect to real trajectories that train the generative model. Jensen-Shannon divergence (JSD) is a well-known similarity metric mainly used for measuring the similarity of two probability distributions [lin1991divergence]. We employ JSD for several human mobility characteristics.
@@
In addition to divergence metrics, we introduce a new metric that measures link trajectory connectivity as a percentage. The connectivity metric quantifies the percentage of trajectories that are fully connected from origin to destination, owing to the RCM that samples the next road link conditioned on the connected road links, thereby capturing the underlying road network topology and connectivity patterns. We employed the connectivity metric in our ablation studies to assess the impact of various components on link trajectory connectivity.
V-D Model Details
MobilityGPT utilizes a minimal version of the standard decoder-only transformer architecture [radford2019language], built upon the minimal GPT implementation from [karpathy_mingpt]. Table I summarizes the important model and training parameters employed in this work. Since the Porto and Beijing datasets exhibit different trajectory lengths and characteristics, we utilize distinct block sizes for training and maximum trajectory lengths for synthetic trajectory generation. However, we employ the same set of hyperparameters across both datasets for all training stages, including the reward model and PPO training.
TABLE I: Model and training parameters utilized in MobilityGPT
@@ V-D Model Details
V-E Comparison with benchmark generative methods
TABLE II: Comparison of the utility metrics with benchmark studies. Bold and underlined results show the best and second-best results, respectively. For all metrics, lower values indicate better performance. While query error relies on normalized error rates for most visited places, other metrics use the Jensen-Shannon distributional divergence statistic.
  Porto-Taxi BJ-Taxi
Random Walk 1.729 0.375 0.739 0.393 0.307 0.105 0.167 0.473 0.187 0.133
MMC 1.409 0.210 0.742 0.393 0.307 0.212 0.167 0.550 0.465 0.291
SeqGAN 0.138 0.116 0.220 0.191 0.326 0.063 0.296 0.162 0.275 0.120
DiffTraj 0.267 0.177 0.126 0.123 0.252 0.074 0.146 0.158 0.123 0.114
Table II summarizes the performance of MobilityGPT and benchmark models. We sampled $5000$ random trajectories from the test set and generated same number of synthetic trajectories. The results show the significant advantage of MobilityGPT over existing state-of-the-art models for trajectory generation. Across both Porto-Taxi and BJ-Taxi, MobilityGPT consistently outperforms all baseline models, achieving the best or second-best performance on multiple utility metrics.
The results in Table II demonstrate that MobilityGPT consistently outperforms baseline models across both the Porto and Beijing datasets and all utility metrics. Rule-based methods such as Random Walk and MMC show relatively high divergence scores, indicating poor alignment with real-world distributions. While deep generative models like SeqGAN and DiffTraj achieve competitive results on some metrics, they lack consistency across all aspects of trajectory modeling. In contrast, MobilityGPT achieves the lowest JS divergence across all categories on both datasets—for instance, reducing OD distribution divergence to $0.114$ (Porto) and $0.099$ (Beijing), outperforming the next best scores of $0.116$ and $0.105$, respectively. This performance reflects the combined impact of MobilityGPT’s autoregressive transformer architecture and its gravity-based sampling strategy, which better captures global mobility patterns. By grounding generation in structural road connectivity and learned spatio-temporal dynamics, MobilityGPT shows strong potential for applications in transportation, urban planning, and mobility analytics.
(a) Distribution of number of points per trajectory for the Porto-Taxi dataset.
(b) Distribution of number of points per trajectory for the BJ-Taxi dataset.
(c) Distribution of inter-point distance for the Porto-Taxi dataset.
(d) Distribution of inter-point distance for the BJ-Taxi dataset.
Figure 5: Comparison of different trajectory generation models on the Beijing and Porto datasets. The plots show the distribution of the number of points per trajectory and the inter-point distance.
The box plots in Figure 5 illustrate the performance of MobilityGPT and several other models across two key metrics: the number of points per trajectory and inter-point distance. Analysis of the Number of Points per Trajectory reveals that MobilityGPT’s performance is consistently strong. On the Porto dataset (Figure 5(a)), its median of $185.00$ is very close to the Baseline’s $176.00$, demonstrating that MobilityGPT effectively replicates the trajectory length characteristics of this dataset. Similarly, for the Beijing dataset (Figure 5(b)), MobilityGPT’s median of $84.00$ is remarkably close to the Baseline’s $89.00$. This indicates that MobilityGPT successfully replicates the fundamental characteristic of trajectory length from the original datasets, a crucial aspect of generating realistic, synthetic mobility data. In contrast, models like TS-TrajGen, MMC, and Random Walk show significantly different distributions, suggesting they generate trajectories that are either excessively long or follow a different statistical pattern than the baseline.
Regarding Inter-point Distance, MobilityGPT also demonstrates a high degree of fidelity to the baseline data. For the Porto dataset (Figure 5(c)), the results are highly aligned, with MobilityGPT’s mean of $20.77$ meters and median of $10.01$ meters nearly identical to the Baseline’s $20.59$ meters and $10.06$ meters, respectively. On the Beijing dataset (Figure 5(d)), MobilityGPT’s mean of $61.98$ meters and median of $26.18$ meters are highly comparable to the Baseline’s mean of $61.20$ meters and median of $25.21$ meters. These results underscore that MobilityGPT not only generates trajectories of the correct length but also accurately captures the realistic step distances between consecutive points. This is a significant finding, as models like DiffTraj and SeqGAN exhibit a much wider variance or different median values, suggesting they may struggle to accurately represent the typical movement patterns found in real-world data. Overall, the consistent alignment of MobilityGPT’s metrics with the Baseline across both datasets confirms its effectiveness in generating high-fidelity, synthetic trajectory data.
(a) Original Trajectories
(b) MobilityGPT
(c) TS-TrajGen
(d) SeqGAN
(e) LSTM-TrajGAN
(f) DiffTraj
V-F Ablation study
As MobilityGPT incorporates three key enhancements to the standard GPT model for capturing human mobility patterns — gravity model, RCM, and RLTF — we conducted ablation experiments to evaluate their individual and combined contributions. Results, presented in Table V-Fa and LABEL:tab:ablation_study_BJb, show that removing the gravity model and RCM leads to noticeable performance declines. The gravity model plays a critical role in modeling trip lengths, offering substantial improvements even when used in isolation. Without the RCM, MobilityGPT struggles to maintain spatial connectivity in the generated trajectories. By conditioning the next-link prediction on actual road connections, RCM effectively enforces geospatial constraints. Furthermore, fine-tuning MobilityGPT with RLTF improves synthetic trajectory quality—particularly trip length—since the preference dataset is optimized for that objective. These results highlight how each component contributes uniquely to different aspects of mobility modeling, and together, they deliver comprehensive improvements across all evaluation metrics.
TABLE III: Ablation study on Porto (a) and Beijing (b) datasets examining the impact of different components: RCM, gravity model, and RLTF. Lower values are better for OD and Trip Length, while higher values are better for Connectivity.